40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Fri Jun 26 16:53:38 2020, gautam, Update, Electronics, RFPD characterization WFS1.pdfWFS2.pdfbuildupMons.pdf
    Reply  Mon Jun 29 15:56:02 2020, gautam, Update, Electronics, RFPD characterization 
    Reply  Sun Jul 5 18:25:42 2020, rana, Update, Electronics, WFS characterization WFS_attenOff.pdf
       Reply  Sun Jul 12 22:40:35 2020, gautam, Update, Electronics, WFS characterization - old SURF report 
       Reply  Sat Aug 29 20:07:48 2020, rana, Update, Electronics, WFS characterization 
          Reply  Sat Aug 29 22:10:09 2020, gautam, Update, Electronics, WFS characterization 
          Reply  Tue Sep 15 17:04:43 2020, gautam, Update, Electronics, DC adaptors delivered 
Message ID: 15450     Entry time: Sun Jul 5 18:25:42 2020     In reply to: 15433     Reply to this: 15472   15547
Author: rana 
Type: Update 
Category: Electronics 
Subject: WFS characterization 

in the lab, checkin on the WFS

Sun Jul 5 18:25:50 2020

I redid Gautam's measurements to get a baseline before changing the head, and my results are very different: To me it looks like the WFS2 quadrants are all OK.

 

Measurement Details:

  1. The whole AG4395 + breadboard Jenne laser is wheeled over next to the SW side of the AP table.
  2. The output of the 1611 goes into channel R of the 4395
  3. I disconnected all the LEMO cables from the head and then plugged a LEMO-BNC cable into the plugs one at a time. The existing LEMO connectors, which take the signals back to the demode board, were all a little loose, so I adjusted them with some pliers (see video).
  4. The Atten = 0 dB for all AG4395 channels
  5. Source drive = 0 dBm. Checked with a -10 dBm drive that there was no change in the observed TFs, so I guess a 0 dBm drive doesn't make things nonlinear.
  6. When I first turned the setup on, the Yellow 'limit' light was ON on the ILX laser current driver, so maybe the modulation wasn't getting to the laser diode as we wish.
  7. did not change any WFS MEDM settings for these measurements. Not sure if any of those buttons work anyway.

I've left the setup as is in case either me or Gautam want to double check. If we're agreed on this response, I'll remove the notches and disable the RF attenuators.

Sun Jul 5 21:42:45 2020

Attachment 1: WFS_attenOff.pdf  404 kB  Uploaded Sun Jul 5 22:12:18 2020  | Hide | Hide all
WFS_attenOff.pdf
ELOG V3.1.3-